1. Determine thedisplacement of a periodic motion given that the time it takes to complete one cycle is 1 second and the amplitude is 1 meter (y=-3.071790 x 10-⁷ m)



2. What is the angular frequency of an oscillating device at a period of 150 ms? (w=2400°/sec or 41.8879 rad/sec)



3. A pendulum is observed to complete 20 full cycles in 58 seconds. Determine the frequency (f = 0.34 cycle/sec)



4. Determine the frequencies of the following periods


a. 30 seconds


b. 3 seconds


c. 30 ms


d. 3000 seconds

1
Expert's answer
2021-11-24T12:00:30-0500

1. The displacement in 1 s is


"y=A\sin(\omega t)=A\sin\bigg(\frac{2\pi}Tt\bigg)=0."

2. We have


"\omega=2\pi/T=41.89\text{ rad/s}."

3. The frequency is


"f=\frac1T=\frac Nt=0.34\text{ Hz}."


4. The frequencies are


"f_a=\frac 1{T_a}=0.033\text{ Hz},\\\space\\
f_b=\frac 1{T_b}=0.33\text{ Hz},\\\space\\
f_c=\frac 1{T_c}=33\text{ Hz},\\\space\\
f_d=\frac 1{T_d}=3.3·10^{-4}\text{ Hz}."